AES-MOM ME/H2O  April - May 2013 Shutdown Work List
LINAC  Open 4/26 6:00am - Close 5/15      
Disconnect Water: solenoids, waveguide, structure in preparation for the removal of components related to the installation for the new PC Gun 31793 Dench RW
Remove and replace existing control enclosure for the L4 water station 31799 Dench RW
Inspect 12 Kw heater wiring on all Linac water stations. Repair heat damaged wire and components as necessary 31794 Dench RW
Temporarily remove shielding wall (LS-02) from between the Linac and the Linac Gallery 31715   ET
Remove the Gun Solenoid magnet, stand and Laser support along with the support grout 31714 Bracs ET
Remove the Solenoid Magnet from the L1AS1 Accelerating Structure 31713 Bracs ET
PAR   Open 4/26 6:00am - Close 5/15          
Inspect Kicker Magnets   DEB CCP    
BOOSTER    Open 4/25 6:00am - Close 5/15      
Replace Magnet hoses     ET    
Hook up Jaski's Magnemt in MM1          
Bldg. 420        
Citranox Griswold cartridges in all RF circulators 31760 Bracs Swetin  
Install new shut off valve in pressure relief line for RF4 water system.     Swetin  
STORAGE RING   Open 4/26 6:00am - Close 5/15        
Go-No-Go 31815 DC/CB ET
Replace BH3 Corrector hoses S15-S22     ET
Sextupole Hoses Replacement in S38     ET
Move the BHBV4 corrector magnet from girder #3 to between girders #3 and #4 in sectors 34 and 1 (Fast Steering Corrector Relocation Project) 31816 MRJ ET
Inspect Klixons and thermal switches at the cabinets as well as a visual inspection for radiation damage in the tunnel 31792 CB/RR RW
S37 Scaper installation 31938 BF RB
Section 6 bakeout in 27 and 28 (disconnect and blow out) 31825 | 31826 GM/DB  
Full Bake-out in Sectors 1 and 2 31796 BF/TN RW
Apply Aluminum tape to the edges of clad absorber sheilding   CB ET
S3A:H4, S5A:H4, S7A:H4,  S9A:H4, S11A:H4, S22A:H4, S24A:H4, S26A:H4, S28A:H4, S30A:H4, S32A:H4, S34A:H4, S36A:H4, S38A:H4, S40A:H4 31937 MRJ RRB
Support Back Scattering BL removal Sector 1 31812 BF ET
Repair leaks in S35B:S1, 11BQ4, S35B:S2, HET Dipole, S32A:S3 32001   RB
Upgrade ID front-end interlocks to Yokogawa in 23ID 31755 GM Swetin
Citranox front-end components: 8ID-FM1,PS1; 9ID-PS1,FM1; 17ID-PS1,FM1; 20ID-FM1,PS1 31896 BF/TN Swetin
Install new control valve for pump system P-SR-02 (Lock-out and drain for contractors) 31751 Bracs/Rice Swetin
Pump oil changes 31757   Swetin    
Pump Filter Changes (See WR for locations) 31721   Swetin    
Replace UV Lights 31635   Swetin    
Install new HX on vacuum cooling skid SRVC-22 and SRVC-24 31759 Bracs/Rice Swetin  
Install heater in SRVC-14 31299   Swetin  
SPX pump valve calibration 31761   Swetin  
Install flow meter in vacuum chamber cooling skid sectors 13/14. 31764   Swetin  
Trouble shoot Flow Meter on Storage Ring East Pump.  31802   Wright  
Replace RTD on vacuum chamber cooling water skids for sectors 11/12 and 13/14. 31933   Swetin  
Replace V-cone flow meter for beamline 4. 31717   Swetin  
A 2HP Grundfos water pump located in the mechanical mezzanine between columns 82 and 83 (Doug Robinson Ticket #545531)        
Cryo Pump Maintenance - Sectors 1, 20, 22, 24 31940 DB RB  
Validate PSS Yokogowa Transmitters. (Sectors) 31817   ET
34ID-B Install Polarizer (For Erika Benda ) 31941   RB
A new X-ray shutter needs to be installed into the 14-ID-A hutch MOM Request      
12BM-A Rebuild Header      
Connect 6BM - Install water headers and flow meters for 6BM-A, and 6BM-B 31762   Swetin
Reroute 23ID-C water lines (Dale Ferguson Ticket #944904)      
Sector 17 Hutch Railing      
Look at Jenias Gearbox (Ticket #232113)      
Shutter covers see (Ticket #792831)      
Re-plumb level meter for make-up tank     RW  
Reduce pump speed-EAA Pump to remain running at 100% Capacity -  Notify Tony Putkammer if this pump status needs to change   Bracs/Rice Swetin  
Tasks in progress        
Completed Task      
Deferred Task          
RSS Component CCWP Required        
** Could be done after tunnel closes        
Task added after start of Shutdown      
Updated: 4/10/2013 - 8:52 AM        
EAA Pump to remain running at 100% Capacity -  Notify Tony Putkammer if this pump status needs to change